OpenStudy (anonymous):

sin15*cos15 ( degrees)

OpenStudy (anonymous):

We have still not learnnt the half angle formulae so how to do it without that

OpenStudy (primeralph):

Do you know double angle?

Parth (parthkohli):

And the double angle formulas? Let\[k = \sin(15)\cos(15)\]But\[2k =2\sin(15)\cos(15) = \sin(30) = \dfrac{1}{2}\]\[\Rightarrow k = \dfrac{1}{4}\]

OpenStudy (anonymous):

i know all of them actually but is there a way to solve it without them becoz in school we havent learnt them

OpenStudy (primeralph):

Do you know trig?

OpenStudy (primeralph):

And geometry?

OpenStudy (anonymous):

are u kidding @primeralph

OpenStudy (primeralph):

I can't read your mind.

OpenStudy (primeralph):